Hi can I have some help concerning functions?
A quadratic function is given.
f(x) = x^2 - 9x
(a) Express the quadratic function in standard form.
f(x) = a(x - h)^2 + k where
a =
h =
k =

Answer by scott8148(6628)

to get the "perfect square" part of the function, divide the x coefficient by 2 and add the square of the result

since adding a quantity to the function would change the function, you must also subtract the same amount

f(x)=x^2-9x __ f(x)=x^2-9x+((-9/2)^2)-((-9/2)^2) __ f(x)=(x-(9/2))^2-((-9/2)^2)

a=1 __ h=9/2 __ k=-((-9/2)^2)